Circular No. 7938

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA IN M31 An independent discovery of one of the apparent novae in M31 reported

on IAUC 7937 has been received from M. Fiaschi, F. Di Mille, and R. Cariolato, Astronomical Observatory ‘G. Colombo’, Padua, Italy. The nova is located at α = 0^{h}43^{m}01^{s}.00, δ = +41^{o}16′20′′.3 (equinox 2000.0), which is 189′′.3 east and 11′′.4 north of the nucleus of M31. The CCD images were taken with a 0.41-m reflector through Hα, R, and B filters, yielding the following magnitudes: July 5.97 UT, R = 17.0; 5.98, V = 17.4; 5.99, B ∼ 18; 11.06, Hα = 16.7; 11.99, R ∼ 18; 12.03, Hα = 16.3. An image of the same field through an Hα filter on Jan. 25 UT showed nothing at this position (limiting mag 17.5).

V2540 OPHIUCHI R. C. Puetter, Center for Astrophysics and Space Science, University

of California at San Diego; D. K. Lynch, S. Mazuk, R. J. Rudy, C. C. Venturini, The Aerospace Corporation; and R. B. Perry, NASA Langley Research Center, report 0.8–2.5-μm spectroscopy using the Lick Observatory 3-m telescope and Aerospace Corporation Near-Infrared Imaging Spectrograph on July 17.3 UT, ∼ 172 days after discovery: “The nova shows a very rich emission line spectrum. Although He I 1.0830 μm is the strongest emission line, low excitation features of O I and Fe II are still present. The O I lines are fluorescently excited, both by the continuum and by H I Lyβ. The so-called one μm iron lines (Fe II 0.9997, 1.0501, 1.0863, and 1.1126 μm) are the strongest we have seen in a nova. The He II lines have not yet appeared nor have the coronal lines. The H I and He I lines have both narrow (FWHM ∼ 900 km/s) and broad (FWHM ∼ 3000 km/s) components.”

Visual and CCD magnitude estimates: Feb. 23.33 UT, 9.3 (A. Amorim, Florianopolis, Brazil); Mar. 8.81, 8.8 (S. Yoshida, Ibaraki, Japan); Apr. 5.80, 8.8 (A. Pearce, Nedlands, W. Australia); June 2.95, 10.4 (R. Y. Shida, São Paulo, Brazil); July 12.13, V = 11.18 ± 0.05 (J.D. West, Mulvane, KS)

SUPERNOVA 2002ec IN NGC 5910 M. M. M. Santangelo, Lucca, Italy, reports that patrol CCD images

of NGC 5910 taken with a 0.51-m reflector in the course of the Monte Agliale Supernova Search on June 15.01 and 26.95 UT show nothing near the position of SN 2002ec (limiting R mag ∼ 20.0).

2002 July 27 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT Carl W. Hergenrother